This is an equation in quadratic form.
I need to find all roots.
What are the steps. I am unclear on how to substitute.
(3/n) - (7/(sqrt n)) - (6) = 0
I think that the answer is 1/9, but I am not sure if I substituted right.

Answer by khwang(438)

Let m = 1/sqrt(n), so 1/n = m^2.
Substitute into (3/n) - (7/(sqrt n)) - (6) = 0 we have
3m^2 - 7m - 6 = 0,
Factor: (3m + 2)(m -3) = 0
So,m = 3 or m =-2/3 (negative, invalid)
Hence, n = 1/m^2 = 1/9...Answer
Check: (3/n) - (7/(sqrt n)) - 6 = 27 -21 -6 = 0 (OK)

Kenny